If I have an isosceles triangle, and the hypotenuse is SQRT(6), what is the length of one leg?
Answer by richwmiller(17219)

If you have an isosceles triangle with a hypotenuse then it is an isosceles right triangle with equal legs.
2a^2=c^2
2a^2=(sqrt(6))^2
2a^2=6
a^2=3
a=sqrt(3)
one leg =sqrt(3)